OpenStudy (anonymous):

The ages of two brothers are in the ratio two to three, but in eight years, the ratio of their ages will be three to four. What is the age of the younger brother? 16 18 24

OpenStudy (anonymous):

i dont know how to do this tho

OpenStudy (nurali):

let x = age of younger bro ley y = age of older bro x/y = 2/3 ----> x = 2y/3 eq.1 x+8 / y+8 = 3/4 eq.2 sub eq.1 into eq.2: (2y/3) + 8 / y+8 = 3/4 (2y+24)/3 / y+8 = 3/4 2y + 24 / 3(y+8) = 3/4 2y+24 / 3y+24 = 3/4 4(2y+24) = 3(3y+24) 8y+96 = 9y+72 24=y x=2y/3=2*24/3=16

OpenStudy (anonymous):

how did you get that so fast

OpenStudy (anonymous):

he went to yahooo awnsers i figured him out

OpenStudy (mathmale):

This is clearly a problem in Algebra. Nurali did very well in starting out by identifying and naming his unknowns: let x = age of younger bro let y = age of older bro Then he wrote two equations showing how these ages (x and y) are related: Right now the ratio of the ages of the brothers is 2 to 3: \[\frac{ x }{ y }=\frac{ 2 }{ 3 }, or 3x=2y\] after cross multiplication.

OpenStudy (mathmale):

Then he moved forward eight years and re-wrote that ratio: \[\frac{ x+8 }{ y+8 }=\frac{ 3 }{ 4 }\] You now have 2 equations in 2 unknowns. The first one, 3x=2y, can be solved for either x or y. Solving for x, x=(2y)/3. Substitute this expression for x into the 2nd equation. I'm not saying this work is easy, but at least y ou're on the right track.